OpenStudy (anonymous):

Simplify, showing your steps. Tell the property you used in each step. 1/2(1*2)+ -1

OpenStudy (mathstudent55):

a * 1 = a (multiplicative identity property) a * 1/a = 1 (multiplicative inverse, or reciprocal) a + (-a) = 0 (additive inverse)

OpenStudy (anonymous):

Thank you

OpenStudy (mathstudent55):

You're welcome.